On September 3, 2021, the Robers Company exchanged equipment with Phifer Corporation. The facts of the exchange are as follows: Robers’ Asset Phifer’s Asset Original cost $ 195,000 $ 215,000 Accumulated depreciation 115,000 123,000 Fair value 97,500 77,500 To equalize the exchange, Phifer paid Robers $20,000 in cash.

Required: Record the exchange for both Robers and Phifer. The exchange has commercial substance for both companies.
Answer #1

In books of Rober's company :

DATE ACCOUNT TITLE DEBIT CREDIT
September 3 ,2021 Cash 20000
Equipment (New) 77500
Accumulated depreciation (old) 115000
Equipment (old) 195000
Gain on exchange of assets 17500

Total Amount Received =Cash +fair market value of asset received

                            = 20000 + 77500

                            = 97500

Book value of asset given =cost -Accumulated depreciation

                           = 195000 - 115000

                            = 80000

Gain on exchange on asset = 97500- 80000 = 17500

B)In Books of Phifer company :

DATE ACCOUNT TITLE DEBIT CREDIT
September 3,2021 Equipment (New) 97500
Accumulated depreciation (old) 123000
Loss on exchange of asset 14500
Equipment (old) 215000
cash 20000

**Fair value of asset received = 97500

Amount paid = 20000+ [215000-123000]Book value of old asset

                 = 20000+ 92000

              = 112000

Loss on exchange of asset = 97500- 112000 = - 14500